You are standing 49 meters from the base of a building. You estimate that the angle of elevation to the top of the 86th floor (the observatory) is 82°. If the total height of the building is another 121 meters above the 86th floor, what is the approximate height of the building? (Round your answers to one decimal place.)

Answer by stanbon(75887)

You are standing 49 meters from the base of a building. You estimate that the angle of elevation to the top of the 86th floor (the observatory) is 82°. If the total height of the building is another 121 meters above the 86th floor, what is the approximate height of the building? (Round your answers to one decimal place.)
---------------
Sketch the picture.
You have a right triangle with base angle = 82 deg, base = 49 meters.
Find the height::
height = base*tan(82)
h = 49*7.115 = 348.65 meters (height to top of the 86th floor)
-----
Total height = 348.65 + 121 = 469.65 meters = 469.7 meters when rounded
---------------
Cheers,
Stan H.
---------------------